The perimeter of a triangle is 30 inches. the first side of the triangle is 1 inch longer than the second side. the second side is 7 inches longer than the third side. what is the length of each side of the triangle?
Answer by fractalier(6550)

Call the sides a, b and c. Then we have
a + b + c = 30
a = b + 1
b = c + 7 or c = b - 7
Now substitute into the first equation and get
b+1 + b + b-7 = 30
3b - 6 = 30
3b = 36
b = 12 so that
a = 13
c = 5